Check whether the function: f: R^2→R, defined by

f(x,y)= x+ysinx has extremum at any point in the domain of f.



1
Expert's answer
2021-04-14T07:36:59-0400

"f(x, y)=x+y\sin{x}"


"f'_y(x, y)=\sin{x}"

derivative equals 0 when "x=\pi*n" , "n=0,\pm1,\pm2,..."


"f'_x(x, y)=1+y\cos{x}"

derivative equals 0 when "y=-\frac{1}{\cos{x}}" .

"y=-1" when "n" is even and "y=1" when "n" is odd.


"f''_{xx}(x, y)=-y\sin{x}"


"A=f''_{xx}(\pi*n,\pm1)=0"


"f''_{yy}(x,y)=0"


"B=f''_{yy}(\pi*n,\pm1)=0"


"f''_{xy}(x, y)=\cos{x}"


"D=f''_{xy}(\pi*n,\pm1)=\pm1"


"AB-D^2=-1<0"


so the function has no extremum